How many grams of solid ammonium bromide should be added to 1.00 L of a 0.293 M ammonia solution to prepare a buffer with a pH of 8.610 ?

grams ammonium bromide = g.
Answer #1

mixture of NH3 and NH4Br act as basic buffer

pOH = pKb + log [NH4Br] / [NH3]

pH =8.610

pOH = 14 - 8.610 = 5.390

pKb of NH3 = 4.74

5.390 = 4.74 + log [NH4Br] /[0.293]

0.65 = log [NH4Br] - log [0.293]

0.65 = log [NH4Br] + 0.53

log [NH4Br] = 0.15

[NH4Br] = 1.32 M

now

Molarity = (W /MW) (1 /V in L)

1.32 = (W / 97.94)(1/1)

W = 129.3 g

mass of NH4Br required = 129.3 g